OpenStudy (anonymous):

how do i solve this? sin3x=-1/2

OpenStudy (anonymous):

solve sin(y)=-1/2 Then solve 3x = y can you do that?

OpenStudy (anonymous):

y = 90?

OpenStudy (anonymous):

y = 90?

OpenStudy (anonymous):

Nope it is -90 or 270

OpenStudy (shubhamsrg):

its not 270, its only -90 see,, you're using inverse here,,got to be in the domain..

OpenStudy (kropot72):

\[\sin ^{-1}(-\frac{1}{2})=-30degrees\]

OpenStudy (shubhamsrg):

ohh yes,,lol..-30

OpenStudy (shubhamsrg):

and not 210

OpenStudy (kropot72):

y = (360 - 30) degrees

OpenStudy (kropot72):

\[y=3x=(360-30)^{o}\] \[x=\frac{(360-30)}{3}degrees\]